# On finding optimal collective variables for complex systems by minimizing the deviation between effective and full dynamics

### v1

@@ 3.2 Properties of effective dynamics
where the first equality follows from the expression of $\bar{p}$ in (47), the second equality follows by applying the identity (42) and the definition of $\widetilde{p}$ in (44), the third equality follows from the invariance of $\pi$ in (1), the fourth equality follows by applying the identity (42) and the definition of $\mu_{z}$ in (39), and finally the last equality follows from the definition of KL-divergence.
Putting (49) and (50) together, we get, for any $\bar{p}\in\Theta^{\xi}$,
$\begin{split}&\mathbb{E}_{x\sim\mu}\Big{(}D_{\mathrm{KL}}\big{(}p(x,\cdot)\,\|\,\bar{p}(x,\cdot)\big{)}\Big{)}\\ =&\int_{\mathbb{R}^{k}}\bigg{[}D_{\mathrm{KL}}\big{(}\widetilde{p}(z,\cdot)\,\|\,g(z,\cdot)\big{)}+D_{\mathrm{KL}}(p_{z}\,\|\,\bar{p}_{z})\bigg{]}Q(z)\,\mathrm{d}{z}\\ &-\int_{\mathbb{R}^{k}}\bigg{[}\int_{\mathbb{R}^{k}}(\widetilde{p}\ln\widetilde{p})(z,\hat{z})\,\mathrm{d}{\hat{z}}+\int_{\Sigma_{z}}\ln p_{z}(x)\mu_{z}(\mathrm{d}{x})\bigg{]}Q(z)\,\mathrm{d}{z}\\ &+\int_{\mathbb{R}^{d}}\bigg{[}\int_{\mathbb{R}^{d}}p(x,y)\ln p(x,y)\,\mathrm{d}{y}\bigg{]}\pi(x)\,\mathrm{d}{x}\,.\end{split}$ (51)
@@
Since $\bar{p}$ is only involved (through $g$ and $\bar{p}_{z}$) in the first integral on the right hand side of (51) and the KL-divergence is nonnegative and it equals to zero if and only if the two densities are identical, we conclude that the objective in (48) is minimized when $\bar{p}$ is given in (47) with $g=\widetilde{p}$ and $\bar{p}_{z}=p_{z}$ for $z\in\mathbb{R}^{k}$. ∎
@@
$\min_{\xi}\min_{\bar{p}\in\Theta^{\xi}}\mathbb{E}_{x\sim\mu}\Big{(}D_{\mathrm{KL}}\big{(}p(x,\cdot)\,\|\,\bar{p}(x,\cdot)\big{)}\Big{)}$ (52)
$\Longleftrightarrow$ $\min_{\xi,g,\bar{p}_{z}}\bigg{[}-\int_{\mathbb{R}^{d}}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)\ln g(\xi(x),\xi(y))\,\mathrm{d}{y}\Big{]}\pi(x)\,\mathrm{d}{x}-\int_{\mathbb{R}^{d}}\ln\bar{p}_{\xi(x)}(x)\pi(x)\,\mathrm{d}{x}\bigg{]}$ (53)
$\Longleftrightarrow$ $\min_{\xi,g}\bigg{[}-\int_{\mathbb{R}^{d}}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)\ln g(\xi(x),\xi(y))\,\mathrm{d}{y}\Big{]}\pi(x)\,\mathrm{d}{x}+\max_{\bar{Q}}\int_{\mathbb{R}^{d}}\ln\bar{Q}(\xi(x))\pi(x)\,\mathrm{d}{x}$
$\hskip 34.14322pt+\frac{1}{2}\int_{\mathbb{R}^{d}}\ln\mathrm{det}(\nabla\xi(x)\nabla\xi(x)^{\top})\pi(x)\,\mathrm{d}{x}\bigg{]}\,,$ (54)
where $g$ is a transition density on $\mathbb{R}^{k}$, $\bar{p}_{z}$ is a probability density on $\Sigma_{z}$, and $\bar{Q}$ is a probability density on $\mathbb{R}^{k}$
@@
The equivalence between (52) and (53) follows by rewriting (51) using (42) and noticing the fact that the last term there is independent of both $\xi$ and $\bar{p}$. By a similar argument, from (51) and using the fact that the divergence $D_{KL}(p_{z}\|\bar{p}_{z})$ vanishes when $\bar{p}_{z}=p_{z}$, we obtain that (52) is equivalent to
$\min_{\xi,g}\bigg{[}-\int_{\mathbb{R}^{d}}\Big{[}\int_{\mathbb{R}^{d}}p(x,y)\ln g(\xi(x),\xi(y))\,\mathrm{d}{y}\Big{]}\,\pi(x)\,\mathrm{d}{x}-\int_{\mathbb{R}^{d}}\ln p_{\xi(x)}(x)\pi(x)\,\mathrm{d}{x}\bigg{]}\,.$ (55)
To show the equivalence between (52) and (54), note that, using the expression of $D_{KL}(Q|\bar{Q})$ where $Q$ is the probability density on $\mathbb{R}^{k}$ defined in (40), we can verify that
$\int_{\mathbb{R}^{d}}\ln Q(\xi(x))\pi(x)\,\mathrm{d}{x}=\max_{\bar{Q}}\int_{\mathbb{R}^{d}}\ln\bar{Q}(\xi(x))\pi(x)\,\mathrm{d}{x}.$
Using the identity above and the expression of the density $p_{z}$ in (85), we can rewrite (55) as (54). This shows the equivalence between (52) and (54). ∎
@@
Theorem 3.1 provides two possible data-driven numerical approaches for jointly learning $\xi$ and the transition density $\widetilde{p}$ of the effective dynamics. (see [3] for alternative objective loss functions involving weighted $L^{1}$ norm of densities).
@@
    $\min_{\xi}\min_{\bar{p}\in\Theta^{\xi}}\Big{[}-\frac{1}{N}\sum_{n=0}^{N-1}\ln\bar{p}(X_{n},X_{n+1})\Big{]}\,.$ (56)
    This recovers the loss function proposed in [51] for learning CV maps based on normalizing flows [23].
@@
    $\begin{split}\min_{\xi}\Big{[}&-\max_{g}\frac{1}{N}\sum_{n=0}^{N-1}\ln g(\xi(X_{n}),\xi(X_{n+1}))+\max_{\bar{Q}}\frac{1}{N}\sum_{n=0}^{N-1}\ln\bar{Q}(\xi(X_{n}))\\ &+\frac{1}{2N}\sum_{n=0}^{N-1}\ln\mathrm{det}(\nabla\xi(X_{n})\nabla\xi(X_{n})^{\top})\Big{]}\,.\end{split}$ (57)
In (56), learning $\bar{p}$ requires to learn the densities $\bar{p}_{z}$ on (high-dimensional) level sets. In contrast, both $g$ and $\bar{Q}$ in (57) are functions in a low-dimensional space, i.e. $\mathbb{R}^{k}$. Hence, the second approach may have less complexity in comparison with the first approach.
@@
$(\widetilde{\mathcal{T}}\widetilde{f}\,)(z)=\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\widetilde{f}(\hat{z})\,\mathrm{d}{\hat{z}}\,,\quad z\in\mathbb{R}^{k}$ (58)
@@
$\langle\widetilde{f},\widetilde{g}\,\rangle_{\widetilde{\mu}}=\int_{\mathbb{R}^{k}}\widetilde{f}(z)\widetilde{g}(z)Q(z)\,\mathrm{d}{z}\,,\quad\widetilde{f},\widetilde{g}\in\widetilde{\mathcal{H}}\,.$ (59)
@@
    For any $\widetilde{g}\in\widetilde{\mathcal{H}}$, we have $(\widetilde{\mathcal{T}}\,\widetilde{g})(z)=\mathbb{E}_{\mu_{z}}(\mathcal{T}g)$ for all $z\in\mathbb{R}^{k}$, where $g=\widetilde{g}\circ\xi$ and $\mathbb{E}_{\mu_{z}}(\cdot)$ denotes the expectation with respect to $\mu_{z}$.
@@
    For any two functions $\widetilde{f},\widetilde{g}\in\widetilde{\mathcal{H}}$, we have $\langle\widetilde{\mathcal{T}}\widetilde{f},\widetilde{g}\rangle_{\widetilde{\mu}}=\langle\mathcal{T}f,g\rangle_{\mu}$, where $f=\widetilde{f}\circ\xi$ and $g=\widetilde{g}\circ\xi$.
@@
Concerning the first claim, given $\widetilde{g}\in\widetilde{\mathcal{H}}$, let us compute
$(\widetilde{\mathcal{T}}\,\widetilde{g})(z)=$ $\int_{\mathbb{R}^{k}}\widetilde{p}(z,\hat{z})\widetilde{g}(\hat{z})\,\mathrm{d}{\hat{z}}$
$=$ $\int_{\mathbb{R}^{k}}\bigg{[}\int_{\Sigma_{z}}\int_{\mathbb{R}^{d}}p(x,y)\delta(\xi(y)-\hat{z})\,\mathrm{d}{y}\,\mu_{z}(\mathrm{d}{x})\bigg{]}\widetilde{g}(\hat{z})\,\mathrm{d}{\hat{z}}$
$=$ $\int_{\Sigma_{z}}\int_{\mathbb{R}^{d}}p(x,y)\widetilde{g}(\xi(y))\,\mathrm{d}{y}\,\mu_{z}(\mathrm{d}{x})$
$=$ $\mathbb{E}_{\mu_{z}}(\mathcal{T}g)\,,$
where $g=\widetilde{g}\circ\xi$, and we have used (58), (44), as well as (42) to derive the above equalities. Concerning the second claim, using (39), (59), the first claim, and the identity (42), we can compute
$\langle\widetilde{\mathcal{T}}\,\widetilde{f},\widetilde{g}\rangle_{\widetilde{\mu}}=$ $\int_{\mathbb{R}^{k}}\big{(}\mathbb{E}_{\mu_{z}}(\mathcal{T}f)\big{)}\widetilde{g}(z)Q(z)\,\mathrm{d}{z}$
$=$ $\int_{\mathbb{R}^{k}}\bigg{[}\frac{1}{Q(z)}\int_{\mathbb{R}^{d}}(\mathcal{T}f)(x)\delta(\xi(x)-z)\pi(x)\,\mathrm{d}{x}\bigg{]}\widetilde{g}(z)Q(z)\,\mathrm{d}{z}$
$=$ $\int_{\mathbb{R}^{d}}(\mathcal{T}f)(x)\widetilde{g}(\xi(x))\pi(x)\,\mathrm{d}{x}$
$=$ $\langle\mathcal{T}f,g\rangle_{\mu}\,,$
where $f=\widetilde{f}\circ\xi$ and $g=\widetilde{g}\circ\xi$.
@@
Suppose that in applications the VAMP-r score is maximized in order to learn the $m$ leading eigenfunctions $(f_{i})_{1\leq i\leq m}$ of $\mathcal{T}$ (see Remark 2.3 and [50, 31]). Consider first the case $r=1$. Often, these eigenfunctions are represented as functions of system’s features, i.e. $f_{i}(x)=\widetilde{f}_{i}(\mathbf{R}(x))$, where $\widetilde{f}_{i}:\mathbb{R}^{K}\rightarrow\mathbb{R}$, integer $K\gg 1$, and $\mathbf{R}=(\mathbf{R}_{1},\dots,\mathbf{R}_{K})^{\top}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K}$ is a map describing $K$ features (i.e. feature map or internal variables) of the system $X_{n}$. Consider this specific CV map $\xi=\mathbf{R}$ and define the quantities in Section 3.1 accordingly. Using the second claim in Lemma 2, we obtain for the VAMP-1 score
$\sum_{i=1}^{m}\langle f_{i},\mathcal{T}f_{i}\rangle_{\mu}=\sum_{i=1}^{m}\langle\widetilde{f}_{i},\widetilde{\mathcal{T}}\widetilde{f}_{i}\rangle_{\widetilde{\mu}}\,,$ (60)
@@
$\mathbb{E}_{\widetilde{\mu}}(\widetilde{f}_{i})=0,$ $\quad\forall\,i\,\in\{1,\dots,m\}\,,$ (61)
$\langle\widetilde{f}_{i},\widetilde{f}_{j}\rangle_{\widetilde{\mu}}=\delta_{ij}\,,$ $\quad\forall\,i,j\,\in\{1,\dots,m\}\,.$
In other words, approximating eigenfunctions of $\mathcal{T}$ by functions of system’s features is equivalent to approximating eigenfunctions of $\widetilde{\mathcal{T}}$ associated to the feature map $\mathbf{R}$. A similar conclusion holds for the case of VAMP-r score, where $r>1$, since we can obtain a similar identity as (60) by considering effective dynamics built with lag-time $r\tau$ and noticing that $\mathcal{T}^{r}$ is the transfer operator corresponding to the lag-time $r\tau$ (see Remark 2.3).
In the last part of this section, we discuss the reversibility of $Z_{n}$ and its connection to the reversibility of $X_{n}$. Denote by $\widetilde{\mathcal{T}}^{*}$ the adjoint of $\widetilde{\mathcal{T}}$ in $\widetilde{\mathcal{H}}$, which satisfies $\langle\widetilde{\mathcal{T}}\widetilde{f},\widetilde{g}\,\rangle_{\widetilde{\mu}}=\langle\widetilde{f},\widetilde{\mathcal{T}}^{*}\widetilde{g}\,\rangle_{\widetilde{\mu}}$ for any $\widetilde{f},\widetilde{g}\in\widetilde{\mathcal{H}}$. The following result gives the concrete expression of $\widetilde{\mathcal{T}}^{*}$.
@@
$(\widetilde{\mathcal{T}}^{*}\widetilde{f}\,)(z)=\int_{\mathbb{R}^{k}}\widetilde{p}^{*}(z,\hat{z})\widetilde{f}(\hat{z})\,\mathrm{d}{\hat{z}}\,,\quad z\in\mathbb{R}^{k}\,,$ (62)
@@
$\widetilde{p}^{*}(z,\hat{z})=\frac{\widetilde{p}(\hat{z},z)Q(\hat{z})}{Q(z)}=\int_{\Sigma_{z}}\int_{\mathbb{R}^{d}}p^{*}(x,y)\delta(\xi(y)-\hat{z})\,\mathrm{d}{y}\,\mu_{z}(\mathrm{d}{x})\,,\quad z,\hat{z}\in\mathbb{R}^{k}\,,$ (63)
where $p^{*}$ is given in (5).

### v2

@@ 3.2 Properties of effective dynamics
where the first equality follows from the co-area formula, the identity (40) and the definition of $\widetilde{p}$ in (43), and the second equality follows from the definition of KL-divergence.
Similarly, for $I_{2}$, we have
$I_{2}=$ $-\int_{\mathbb{R}^{d}}\ln g_{\xi(x)}(x)\mu(\mathrm{d}{x})$ (53)
$=$ $-\int_{\mathbb{R}^{k}}\left(\int_{\Sigma_{z}}\ln g_{z}(x)\mu_{z}(\mathrm{d}{x})\right)\widetilde{\mu}(\mathrm{d}{z})$
$=$ $-\int_{\mathbb{R}^{k}}\left(\int_{\Sigma_{z}}\ln g_{z}(x)p_{z}(x)\left[\mathrm{det}(\nabla\xi(x)^{\top}\nabla\xi(x))\right]^{-\frac{1}{2}}\sigma_{\Sigma_{z}}(\mathrm{d}{x})\right)\widetilde{\mu}(\mathrm{d}{z})$
$=$ $\int_{\mathbb{R}^{k}}D_{\mathrm{KL}}(p_{z}\,\|\,g_{z})\widetilde{\mu}(\mathrm{d}{z})-\int_{\mathbb{R}^{k}}\bigg{(}\int_{\Sigma_{z}}\ln p_{z}(x)\mu_{z}(\mathrm{d}{x})\bigg{)}\widetilde{\mu}(\mathrm{d}{z})$
$=$ $\int_{\mathbb{R}^{k}}D_{\mathrm{KL}}(p_{z}\,\|\,g_{z})\widetilde{\mu}(\mathrm{d}{z})-\int_{\mathbb{R}^{d}}\ln p_{\xi(x)}(x)\mu(\mathrm{d}{x})\,.$
Hence, we obtain the equalities in (49) by combining (50)–(53).
@@
Since the KL-divergence is nonnegative and it equals to zero if and only if the two densities are identical, we conclude from the last equality in (49) that the objective in (47) is minimized when $g$ is given in (48) with $\widetilde{g}=\widetilde{p}$ and $g_{z}=p_{z}$ for $z\in\mathbb{R}^{k}$. ∎
@@
$\min_{\xi\in\Xi}\min_{g\in\Theta_{\xi}}\mathbb{E}_{x\sim\mu}\Big{(}D_{\mathrm{KL}}\big{(}p(x,\cdot)\,\|\,g(x,\cdot)\big{)}\Big{)}$ (54)
$\Longleftrightarrow$ $\min_{\begin{subarray}{c}\xi\in\Xi,\widetilde{g}\in\widetilde{\mathcal{G}},\\ (g_{z})\in\mathcal{G}_{\xi,\Sigma}\end{subarray}}\bigg{[}-\int_{\mathbb{R}^{d}}\Big{(}\int_{\mathbb{R}^{d}}p(x,y)\ln\widetilde{g}(\xi(x),\xi(y))\,\mathrm{d}{y}\Big{)}\mu(\mathrm{d}{x})-\int_{\mathbb{R}^{d}}\ln g_{\xi(x)}(x)\mu(\mathrm{d}{x})\bigg{]}$ (55)
$\Longleftrightarrow$ $\min_{\xi\in\Xi,\widetilde{g}\in\widetilde{\mathcal{G}}}\bigg{[}-\int_{\mathbb{R}^{d}}\Big{(}\int_{\mathbb{R}^{d}}p(x,y)\ln\widetilde{g}(\xi(x),\xi(y))\,\mathrm{d}{y}\Big{)}\mu(\mathrm{d}{x})+\max_{\widetilde{f}\in\widetilde{\mathcal{S}}}\int_{\mathbb{R}^{d}}\ln\widetilde{f}(\xi(x))\mu(\mathrm{d}{x})\bigg{]}\,,$ (56)
where $\Xi$ is the space of CV maps in (36), $\Theta_{\xi}$ is the space of transition densities defined in (46), and $\widetilde{\mathcal{G}}$, $\widetilde{\mathcal{S}}$, $\mathcal{G}_{\xi,\Sigma}$ are the spaces of densities defined in (45).
@@
We apply the equalities in (49). The equivalence between (54) and (55) follows directly from the first equality in (49), noticing that the second term on the first line of (49) is independent of both $\xi$ and $g$. Next, we show the equivalence between (54) and (56). Note that the second equality in (49) (since the divergence $D_{KL}(p_{z}\|g_{z})$ is nonnegative and it vanishes when $g_{z}=p_{z}$) implies that (54) is equivalent to
$\min_{\xi\in\Xi,\widetilde{g}\in\widetilde{\mathcal{G}}}\bigg{[}-\int_{\mathbb{R}^{d}}\Big{(}\int_{\mathbb{R}^{d}}p(x,y)\ln\widetilde{g}(\xi(x),\xi(y))\,\mathrm{d}{y}\Big{)}\,\mu(\mathrm{d}{x})-\int_{\mathbb{R}^{d}}\ln p_{\xi(x)}(x)\mu(\mathrm{d}{x})\bigg{]}\,.$ (57)
Using the expression of the density $p_{z}$ in (38), we obtain
$\begin{split}&-\int_{\mathbb{R}^{d}}\ln p_{\xi(x)}(x)\mu(\mathrm{d}{x})+\int_{\mathbb{R}^{d}}\ln\pi(x)\mu(\mathrm{d}{x})\\ =&\int_{\mathbb{R}^{d}}\ln\widetilde{\pi}(\xi(x))\mu(\mathrm{d}{x})\\ =&\max_{\widetilde{f}\in\widetilde{\mathcal{S}}}\int_{\mathbb{R}^{d}}\ln\widetilde{f}(\xi(x))\mu(\mathrm{d}{x})\end{split}$ (58)
where $\widetilde{\pi}$ is the probability density on $\mathbb{R}^{k}$ defined in (39), and the second equality can be verified using co-area formula and the fact that $D_{KL}(\widetilde{\pi}\|\widetilde{f})$ is nonnegative for any probability density $\widetilde{f}\in\widetilde{\mathcal{S}}$.
The equivalence between (54) and (56) follows after we substituting (58) into (57) and noticing that the second term on the first line of (58) is independent of both $\xi$ and $\widetilde{g}$. ∎
@@
Theorem 3.1 provides two possible data-driven numerical approaches for jointly learning $\xi$ and the transition density $\widetilde{p}$ of the effective dynamics. (see [7, Section 4] for alternative objective loss functions involving weighted $L^{1}$ norm of densities).
@@
    $\min_{\xi\in\Xi}\min_{g\in\Theta_{\xi}}\Big{[}-\frac{1}{N}\sum_{n=0}^{N-1}\ln g(X_{n},X_{n+1})\Big{]}\,.$ (59)
    This is close to the loss function proposed in [62] for learning CV maps based on normalizing flows [30].
@@
    $\begin{split}\min_{\xi\in\Xi}\Big{[}&-\max_{\widetilde{g}\in\widetilde{\mathcal{G}}}\frac{1}{N}\sum_{n=0}^{N-1}\ln\widetilde{g}(\xi(X_{n}),\xi(X_{n+1}))+\max_{\widetilde{f}\in\widetilde{\mathcal{S}}}\frac{1}{N}\sum_{n=0}^{N-1}\ln\widetilde{f}(\xi(X_{n}))\Big{]}\,.\end{split}$ (60)
In (59), learning $g$ requires to learn the densities $(g_{z})_{z\in\mathbb{R}^{k}}$ on (high-dimensional) level sets. In contrast, both $\widetilde{g}$ and $\widetilde{f}$ in (60) are functions in the low-dimensional space $\mathbb{R}^{k}$. Hence, we expect the second approach to have less complexity in comparison with the first approach.
@@
$(\widetilde{\mathcal{T}}\widetilde{f}\,)(z)=\int_{\mathbb{R}^{k}}\widetilde{p}(z,w)\widetilde{f}(w)\,\mathrm{d}{w}\,,\quad z\in\mathbb{R}^{k}$ (61)
@@
$\langle\widetilde{f},\widetilde{h}\,\rangle_{\widetilde{\mu}}=\int_{\mathbb{R}^{k}}\widetilde{f}(z)\widetilde{h}(z)\widetilde{\mu}(\mathrm{d}{z})\,,\quad\widetilde{f},\widetilde{h}\in\widetilde{\mathcal{H}}\,.$ (62)
@@
    For any $\widetilde{f}\in\widetilde{\mathcal{H}}$, we have $(\widetilde{\mathcal{T}}\,\widetilde{f})(z)=\mathbb{E}_{\mu_{z}}(\mathcal{T}f)$ for all $z\in\mathbb{R}^{k}$, where $f=\widetilde{f}\circ\xi$.
@@
    For any two functions $\widetilde{f},\widetilde{h}\in\widetilde{\mathcal{H}}$, we have $\langle\widetilde{\mathcal{T}}\widetilde{f},\widetilde{h}\rangle_{\widetilde{\mu}}=\langle\mathcal{T}f,h\rangle_{\mu}$, where $f=\widetilde{f}\circ\xi$ and $h=\widetilde{h}\circ\xi$.
@@
Concerning the first claim, given $\widetilde{f}\in\widetilde{\mathcal{H}}$, let us compute
$(\widetilde{\mathcal{T}}\,\widetilde{f})(z)=$ $\int_{\mathbb{R}^{k}}\widetilde{p}(z,w)\widetilde{f}(w)\,\mathrm{d}{w}$
$=$ $\int_{\mathbb{R}^{k}}\left[\widetilde{\pi}(w)\int_{\Sigma_{z}}\left(\int_{\Sigma_{w}}\frac{p(x,y)}{\pi(y)}\mu_{w}(\mathrm{d}{y})\right)\,\mu_{z}(\mathrm{d}{x})\right]\widetilde{f}(w)\,\mathrm{d}{w}$
$=$ $\int_{\Sigma_{z}}\left[\int_{\mathbb{R}^{k}}\left(\int_{\Sigma_{w}}\frac{p(x,y)}{\pi(y)}\mu_{w}(\mathrm{d}{y})\right)\,\widetilde{f}(w)\,\widetilde{\mu}(\mathrm{d}{w})\right]\mu_{z}(\mathrm{d}{x})$
$=$ $\int_{\Sigma_{z}}\left(\int_{\mathbb{R}^{d}}\frac{p(x,y)}{\pi(y)}\widetilde{f}(\xi(y))\mu(\mathrm{d}{y})\right)\mu_{z}(\mathrm{d}{x})$
$=$ $\int_{\Sigma_{z}}\left(\int_{\mathbb{R}^{d}}p(x,y)\widetilde{f}(\xi(y))\mathrm{d}{y}\right)\mu_{z}(\mathrm{d}{x})$
$=$ $\mathbb{E}_{\mu_{z}}(\mathcal{T}f)\,,$
where $f=\widetilde{f}\circ\xi$, and we have used (61), the second equality in (43), as well as (40) to derive the above equalities. Concerning the second claim, using (38), (62), the first claim, and the identity (40), we can compute
$\langle\widetilde{\mathcal{T}}\,\widetilde{f},\widetilde{h}\rangle_{\widetilde{\mu}}=$ $\int_{\mathbb{R}^{k}}\big{(}\mathbb{E}_{\mu_{z}}(\mathcal{T}f)\big{)}\widetilde{h}(z)\widetilde{\mu}(\mathrm{d}{z})$
$=$ $\int_{\mathbb{R}^{k}}\left(\int_{\Sigma_{z}}(\mathcal{T}f)(x)\mu_{z}(\mathrm{d}{x})\right)\widetilde{h}(z)\widetilde{\mu}(\mathrm{d}{z})$
$=$ $\int_{\mathbb{R}^{d}}(\mathcal{T}f)(x)\widetilde{h}(\xi(x))\mu(\mathrm{d}{x})$
$=$ $\langle\mathcal{T}f,h\rangle_{\mu}\,,$
where $f=\widetilde{f}\circ\xi$ and $h=\widetilde{h}\circ\xi$.
@@
Suppose that in applications the VAMP-r score is maximized in order to learn the $m$ leading eigenfunctions $(f_{i})_{1\leq i\leq m}$ of $\mathcal{T}$ (see Remark 2.3 and [61, 40]). Often, these eigenfunctions are represented as functions of system’s features, i.e. $f_{i}(x)=\widetilde{f}_{i}(\mathbf{R}(x))$, where $\widetilde{f}_{i}:\mathbb{R}^{K}\rightarrow\mathbb{R}$, integer $K\gg 1$, and $\mathbf{R}=(\mathbf{R}_{1},\dots,\mathbf{R}_{K})^{\top}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K}$ is a map describing $K$ features (e.g. internal variables) of the system $X_{n}$. Consider this specific CV map $\xi=\mathbf{R}$ and define the quantities in Section 3.1 accordingly. Using the second claim in Lemma 2, we obtain for the VAMP-r score
$\sum_{i=1}^{m}\langle f_{i},\mathcal{T}f_{i}\rangle_{\mu}^{r}=\sum_{i=1}^{m}\langle\widetilde{f}_{i},\widetilde{\mathcal{T}}\widetilde{f}_{i}\rangle_{\widetilde{\mu}}^{r}\,,$ (63)
@@
$\mathbb{E}_{\widetilde{\mu}}(\widetilde{f}_{i})=0,\quad\langle\widetilde{f}_{i},\widetilde{f}_{j}\rangle_{\widetilde{\mu}}=\delta_{ij}\,,\quad\forall\,i,j\,\in\{1,\dots,m\}\,.$ (64)
In other words, approximating eigenfunctions of $\mathcal{T}$ by maximizing VAMP-r score among functions of system’s features is equivalent to approximating eigenfunctions of $\widetilde{\mathcal{T}}$ associated to the feature map $\mathbf{R}$.
A similar remark can be made on the relevance of effective dynamics to numerical algorithms for computing committor based on solving the variational problem in (35). In the last part of this section, we discuss the reversibility of $Z_{n}$ and its connection to the reversibility of $X_{n}$. Denote by $\widetilde{\mathcal{T}}^{*}$ the adjoint of $\widetilde{\mathcal{T}}$ in $\widetilde{\mathcal{H}}$, which satisfies $\langle\widetilde{\mathcal{T}}\widetilde{f},\widetilde{h}\,\rangle_{\widetilde{\mu}}=\langle\widetilde{f},\widetilde{\mathcal{T}}^{*}\widetilde{h}\,\rangle_{\widetilde{\mu}}$ for any $\widetilde{f},\widetilde{h}\in\widetilde{\mathcal{H}}$. The following result gives the concrete expression of $\widetilde{\mathcal{T}}^{*}$.
@@
$(\widetilde{\mathcal{T}}^{*}\widetilde{f}\,)(z)=\int_{\mathbb{R}^{k}}\widetilde{p}^{*}(z,w)\widetilde{f}(w)\,\mathrm{d}{w}\,,\quad z\in\mathbb{R}^{k}\,,$ (65)
@@
$\widetilde{p}^{*}(z,w)=\frac{\widetilde{p}(w,z)\widetilde{\pi}(w)}{\widetilde{\pi}(z)}=$ $\int_{\Sigma_{z}}\left(\int_{\Sigma_{w}}p^{*}(x,y)[\mathrm{det}(\nabla\xi(y)^{\top}\nabla\xi(y))]^{-\frac{1}{2}}\sigma_{\Sigma_{w}}(\mathrm{d}{y})\right)\mu_{z}(\mathrm{d}{x})$ (66)
$=$ $\widetilde{\pi}(w)\int_{\Sigma_{z}}\left(\int_{\Sigma_{w}}\frac{p^{*}(x,y)}{\pi(y)}\mu_{w}(\mathrm{d}{y})\right)\mu_{z}(\mathrm{d}{x})\,,$
for $z,w\in\mathbb{R}^{k}$, where $p^{*}$ is given in (5).